Below a laser shines on a pair of vertical slits G
Answer #1

\lambda = ax/D = 0.31mm*0.511/12.3 = 12878 nm ( May be incorrect)

or \lambda = dy/D = 0.02*0.00031/12.3 = 504 nm (falls in the range)

Here, \lambda is the wavelength , a=d is the aperture, x is the size of proection, D is the distance to the screen ,

y is the size on the screen (maxima)